# Fine-Tuning Discrete Diffusion Models via Reward Optimization with Applications to DNA and Protein Design

### v1

@@ C.1 Straight-Through Gumbel Softmax
C.2 Simplified Formula of $g(\theta)$
@@
C.3 Schedule of Gumbel Softmax Temperature
@@
Appendix D Experimental Details and Additional Results
In this section, we first provide a more detailed explanation of the baseline. We then discuss additional results and present more detailed settings for both regulatory DNA sequence design and protein sequence design.
D.1 Baselines
@@
      The predictor $p(r|x,t)$ is estimated using the posterior mean approach (Chung et al., 2022), where the pretrained model is first utilized to estimate the clean data from the noisy input $x_{t}$, and then the reward oracle is applied to the predicted clean sequence. We remark that the above Taylor approximation doesn’t have formal theoretical guarantees, considering that $x$ is discrete. This could be a reason why it does not work well in the case of protein-inverse folding in Section 6.3.
@@
  • Classifier-free Guidance (CFG) (Ho and Salimans, 2022). Unlike guidance-based methods, CFG trains a conditional generative model from scratch and does not rely on the pretrained model. To generate sequences $x$ with desired properties $r(x)$, CFG incorporates $r(x)$ as an additional input to the diffusion model and generates samples conditioning on high $r(x)$ values. Specifically, binary labels of $r(x)$ are constructed according to the $95\%$ quantile, and sampling is done conditioned on the label corresponding to high values of $r(x)$.
@@
D.2 Regulatory DNA Sequence Design
@@
Fine-tuning Setup. We utilize the pretrained masked discrete diffusion model and the fine-tuning oracle described above for fine-tuning. During DRAKES’s stage 1 data collection, sequences are generated from the pretrained model over 128 steps. We set $\alpha=0.001$ to govern the strength of the KL regularization and truncate the backpropagation at step 50. The model is fine-tuned with 128 samples as a batch (32 samples per iteration, with gradient accumulated over 4 iterations) for 1000 steps. For DRAKES w/o KL, we follow the same setup, but set $\alpha$ to zero. For evaluation, we generate 640 sequences per method (with batch size of 64 over 10 batches) for each random seed. We report the mean and standard deviation of model performance across 3 random seeds.
@@
D.3 Protein Inverse Folding
@@
Finetuning Setup. We utilize the pretrained inverse folding model and the fine-tuning oracle described above for fine-tuning. During DRAKES’s stage 1 data collection, we generate sequences from the pretrained model over 50 steps. We set $\alpha=0.0003$ and truncate the backprogagtion at step 25. The model is finetuned with proteins in Megascale training set with batch size 128 (16 samples per iteration, with gradient accumulated over 8 iterations) for 100 epochs. For DRAKES w/o KL, we follow the same setup, but set $\alpha$ to zero. The model is evaluated on Megascale test set, where we generate 128 sequences conditioned on each protein structure for every method (with batch size of 16 over 8 batches) and each random seed. We report the mean and standard deviation of model performance across 3 random seeds.

### v2

@@ E.1 Straight-Through Gumbel Softmax
E.2 Simplified Formula of $g(\theta)$
@@
E.3 Schedule of Gumbel Softmax Temperature
@@
Appendix F Experimental Details and Additional Results
In this section, we first provide a more detailed explanation of the baseline. We then discuss additional results and present more detailed settings for both regulatory DNA sequence design and protein sequence design. Finally, we present results of DRAKES in the natural language domain for toxicity-controlled text generation. We use 1 NVIDIA A100 80GB GPU for all experiments.
F.1 Baselines
@@
      The predictor $p(r|x,t)$ is estimated using the posterior mean approach (Chung et al., 2022a), where the pretrained model is first utilized to estimate the clean data from the noisy input $x_{t}$, and then the reward oracle is applied to the predicted clean sequence. We remark that the above Taylor approximation doesn’t have formal theoretical guarantees, considering that $x$ is discrete. This could be a reason why it does not work well in the case of protein-inverse folding in Section 6.3.
@@
  • Classifier-free Guidance (CFG) (Ho & Salimans, 2022). Unlike guidance-based methods, CFG trains a conditional generative model from scratch and does not rely on the pretrained model. To generate sequences $x$ with desired properties $r(x)$, CFG incorporates $r(x)$ as an additional input to the diffusion model and generates samples conditioning on high $r(x)$ values. Specifically, binary labels of $r(x)$ are constructed according to the $95\%$ quantile, and sampling is done conditioned on the label corresponding to high values of $r(x)$.
@@
F.2 Regulatory DNA Sequence Design
@@
Fine-tuning Setup. We utilize the pretrained masked discrete diffusion model and the fine-tuning oracle described above for fine-tuning. During DRAKES’s stage 1 data collection, sequences are generated from the model over 128 steps. We set $\alpha=0.001$ to govern the strength of the KL regularization and truncate the backpropagation at step 50. The base Gumbel Softmax temperature is set to 1.0. The model is fine-tuned with 128 samples as a batch (32 samples per iteration, with gradient accumulated over 4 iterations) for 1000 steps. For DRAKES w/o KL, we follow the same setup, but set $\alpha$ to zero. For evaluation, we generate 640 sequences per method (with a batch size of 64 over 10 batches) for each random seed. We report the mean and standard deviation of model performance across 3 random seeds.
@@
Ablation Study on Gradient Truncation Number. To show the impact of the gradient truncation module, we conduct an ablation study on the gradient truncation number. As shown in Table 4, when the truncation number is small, the model cannot effectively update the early stage of the sampling process, leading to suboptimal performance. Meanwhile, when the truncation number is large, the memory cost of the model fine-tuning increases, and the optimization is harder due to the gradient accumulation through the long trajectory. Therefore, an intermediate level of gradient truncation leads to the best performance.
Table 4: Ablation study on gradient truncation number on regulatory DNA sequence design.
Truncation Number Pred-Activity (median) $\uparrow$ ATAC-Acc $\uparrow$ (%) 3-mer Corr $\uparrow$ JASPAR Corr $\uparrow$ App-Log-Lik (median) $\uparrow$ 25 6.17 95.9 0.569 0.798 -278 50 5.61 92.5 0.887 0.911 -264 75 4.61 52.8 0.238 0.644 -268
Ablation Study on Gumbel Softmax Temperature Schedule. We conduct an ablation study with different temperature schedules of Gumbel Softmax, i.e. a linear schedule and a constant schedule. As shown in Table 5, both schedules achieve similar performance, indicating the robustness of our method. In the performance reported in Table 1, we utilize the linear schedule.
Table 5: Ablation study on Gumbel Softmax temperature schedule on regulatory DNA sequence design.
Temperature Schedule Pred-Activity (median) $\uparrow$ ATAC-Acc $\uparrow$ (%) 3-mer Corr $\uparrow$ JASPAR Corr $\uparrow$ App-Log-Lik (median) $\uparrow$ Constant 5.89 91.6 0.852 0.914 -258 Linear 5.61 92.5 0.887 0.911 -264
F.3 Protein Inverse Folding
@@
Finetuning Setup. We utilize the pretrained inverse folding model and the fine-tuning oracle described above for fine-tuning. During DRAKES’s stage 1 data collection, we generate sequences from the model over 50 steps. We set $\alpha=0.0003$ and truncate the backpropagation at step 25. The base Gumbel Softmax temperature is set to 0.5. The model is finetuned with proteins in Megascale training set with batch size 128 (16 samples per iteration, with gradient accumulated over 8 iterations) for 100 epochs. For DRAKES w/o KL, we follow the same setup, but set $\alpha$ to zero. The model is evaluated on Megascale test set, where we generate 128 sequences conditioned on each protein structure for every method (with a batch size of 16 over 8 batches) and each random seed. We report the mean and standard deviation of model performance across 3 random seeds.
@@
Ablation Study on Gradient Truncation Number. To show the impact of the gradient truncation module, we conduct an ablation study on the gradient truncation number. The results are shown in Table 10. While the model performance is robust with different truncation numbers, an intermediate level of gradient truncation leads to the best performance.
Table 10: Ablation study on gradient truncation number on inverse protein folding.
Truncation Number Pred-ddG (median) $\uparrow$ %(ddG$>0$) (%) $\uparrow$ scRMSD (median) $\downarrow$ %(scRMSD$<2$)(%) $\uparrow$ Success Rate (%) $\uparrow$ 15 0.977 83.2 0.852 92.4 76.0 25 1.095 86.4 0.918 91.8 78.6 35 1.033 84.8 0.868 92.7 77.7
Ablation Study on Gumbel Softmax Temperature Schedule. We conduct an ablation study with different temperature schedules of Gumbel Softmax, i.e. a linear schedule and a constant schedule. As shown in Table 11, both schedules achieve similar performance, indicating the robustness of our method. In the performance reported in Table 2, we utilize the linear schedule.
Table 11: Ablation study on Gumbel Softmax temperature schedule on inverse protein folding.
Temperature Schedule Pred-ddG (median) $\uparrow$ %(ddG$>0$) (%) $\uparrow$ scRMSD (median) $\downarrow$ %(scRMSD$<2$)(%) $\uparrow$ Success Rate (%) $\uparrow$ Constant 1.178 86.5 0.897 93.2 80.3 Linear 1.095 86.4 0.918 91.8 78.6
Results with pLDDT. In addition to scRMSD, we further measure the self-consistency of the generated sequences using pLDDT. The results are shown in Table 12. Following the common practice as in Widatalla et al. (2024), we utilize 80 as the cutoff threshold and define a success generation as having positive predicted stability (i.e., Pred-ddG$>$0) and confident ESMFold-predicted structure (i.e., pLDDT $>$ 80). The results align well with those in Table 2 using scRMSD. DRAKES significantly outperforms all baseline methods in terms of the overall success rate.
Table 12: Model performance on inverse protein folding. DRAKES generates protein sequences that have both high stability and high confidence in ESMFold predicted structures, outperforming baselines in the overall success rate. We report the mean across 3 random seeds, with standard deviations in parentheses.
Method Pred-ddG (median) $\uparrow$ %(ddG$>0$) (%) $\uparrow$ pLDDT (median) $\uparrow$ %(pLDDT$>80$)(%) $\uparrow$ Success Rate (%) $\uparrow$ Pretrained -0.544(0.037) 36.6(1.0) 87.9(0.0) 87.5(0.1) 33.8(0.8) CG -0.561(0.045) 36.9(1.1) 87.9(0.0) 87.5(0.2) 34.3(0.5) SMC 0.659(0.044) 68.5(3.1) 88.3(0.1) 84.1(1.0) 53.7(4.3) TDS 0.674(0.086) 68.2(2.4) 88.3(0.1) 85.6(0.7) 54.4(2.8) CFG -1.186(0.035) 11.0(0.4) 65.7(0.0) 7.8(0.1) 0.0(0.0) DRAKES w/o KL 1.108(0.004) 100.0(0.0) 73.4(0.2) 40.4(0.2) 40.4(0.2) DRAKES 1.095(0.026) 86.4(0.2) 87.0(0.0) 85.6(0.7) 72.3(0.8)
Diversity of Generated Sequences. We evaluate the diversity of the protein sequences generated by different methods using the average sequence entropy of each protein backbone in the test set. The results are shown in Table 13. DRAKES achieves a comparable level of diversity as the pretrained model, significantly outperforming SMC and TDS. This further justifies the efficacy of DRAKES to achieve the optimization objective of generating stable sequences, while maintaining high diversity. Among all baselines, DRAKES is the only method that have both a high success rate and high diversity.
Table 13: Average sequence entropy of the generated sequences on protein inverse folding.
Method Sequence Entropy $\uparrow$ Pretrained 34.7(0.2) CG 34.6(0.1) SMC 24.9(1.2) TDS 24.9(0.5) CFG 8.4(0.1) DRAKES w/o KL 25.7(0.1) DRAKES 33.3(0.2)
@@
F.4 Toxicity-Controlled Text Generation
To demonstrate the general applicability of DRAKES beyond its use in the biological domain, we conduct experiments in toxicity-controlled text generation. The corresponding details and results are presented below.
Reward Oracle. We follow the setting in Zhao et al. (2024), using a toxicity prediction model trained by Corrêa (2023) as the reward oracle, which is a fine-tuned RoBERTa model with 125M parameters that can be used to score the toxicity of a sentence.
Pretrained Model. We utilize the pretrained masked diffusion language model from Sahoo et al. (2024) with 130M non-embedding parameters, which is trained on the OpenWebText corpus (Gokaslan et al., 2019) for one million steps and involves the processing of 33 billion tokens.
Fine-tuning Setup. Following (Zhao et al., 2024), we generate 20 output tokens. During DRAKES’s stage 1 data collection, sequences are generated from the model over 20 diffusion steps. We set $\alpha=0.0001$ and truncate the backpropagation at step 5. The base Gumbel Softmax temperature is set to 1.0. The model is fine-tuned with batch size 128 (32 samples per iteration, with gradient accumulated over 4 iterations) for 100 steps. For DRAKES w/o KL, we follow the same setup, but set $\alpha$ to zero. For evaluation, we generation 320 sequences per method (with a batch size of 32 over 10 batches).
Results. As shown in Table 14, DRAKES generates sequences with the highest reward (i.e., low toxicity) compared to all baseline models, measured by the mean and median Pred-Score. This demonstrates the effectiveness of DRAKES in the natural language domain.
Table 14: Model performance on toxicity controlled text generation.
Method Pred-Score (mean) $\uparrow$ Pred-Score (median) $\uparrow$ Pretrained 6.996 8.003 CG 8.488 9.009 SMC 9.286 9.367 TDS 9.817 9.882 DRAKES 10.044 10.049